# Breakdown of chiral anomaly and emergent phases in Weyl semimetals under orbital magnetic fields

### v1

@@ I Introduction
Weyl semimetals (WSMs) [Murakami_2007, Wan_Savrasov_2011, Lv_Ding_2015a, Lv_Ding_2015b, Xu_Hasan_2015a, Xu_Hasan_2015b, Lu_Soljacic_2015] represent a robust gapless topological phase of quantum matter in three dimensions. As long as spatial translation symmetry is maintained, the only way to eliminate this phase is by bringing pairs of Weyl nodes (WNs) with opposite chiralities together, allowing them to annihilate each other. WSMs exhibit a variety of exotic phenomena – the chiral anomaly [Nielsen_Ninomiya_1983, Aji_2012, Zyuzin_Burkov_2012], chiral anomaly-induced negative magnetoresistance [Son_Spivak_2013, Gorbar_Miransky_2014, Burkov_2015, Das_Axialanomaly_2015, Huang_2015, Zhang_Hasan_2016, Li_Das_2016, Lu_Shun_2017], planar Hall effect [Nandy_Tewari_2017, Li_Shen_2018, Shama_Singh_2020, Li_Yao_2023, Wei_Weng_2023], Fermi-arc-mediated quantum oscillations, and three-dimensional quantum Hall effect [Potter_Vishwanath_2014, Zhang_Vishwanath_2016, Moll_Analytis_2016, Wang_Xie_2017, Zhang_Xiu_2019, Li_Xie_2020, Chang_Xing_2021, Ma_Sheng_2021, Chang_Sheng_2022, Zhang_Naoto_2022] – all of which typically rely on the presence of an external orbital magnetic field. However, when the magnetic field strength is sufficiently high, it can couple WNs of opposite chiralities and open a finite gap in the spectrum, thereby destabilizing the WSM state [Goerbig_2009, Zhang_Jia_2017, Ramshaw_McDonald_2018, Chan_Lee_2017, Kim_Park_2017, Saykin_Rodinov_2018, Bednik_Syzranov_2020, Abdulla_Murthy_2022, Abdulla_2024].
@@
An external magnetic field that is not aligned with the separation direction of WNs can couple nodes of opposite chirality, resulting in the opening of a gap in the bulk electronic spectrum. This phenomenon has been the subject of several theoretical investigations based on low-energy continuum models. Notably, Refs. Saykin_Rodinov_2018, Bednik_Syzranov_2020 demonstrated that the application of an orbital magnetic field immediately induces a gap, which is exponentially small, scaling as $\sim\exp[-(Ql_{B})^{2}]$, where $Q$ represents the momentum-space separation between the Weyl nodes and $l_{B}=\sqrt{\hbar/(eB)}$ is the magnetic length associated with the applied field. In these models, the gap increases monotonically as the magnetic field strength is raised. Furthermore, Ref. Devakul_Parameswaran_2021 revealed that in systems where the Weyl fermion velocities exhibit anisotropy, the induced gap can display oscillatory behavior as a function of the magnetic field strength, adding further complexity to the field-induced modification of the electronic structure.
@@
It is important to note that these results, derived from continuum approximations, are strictly valid within a specific regime of parameters: when the separation between the Weyl nodes ($Q$) is much smaller than the linear dimensions of the Brillouin zone (BZ), and when the magnetic length $l_{B}$ is considerably larger than the lattice constant $a$, so that lattice effects can be safely ignored. However, these conditions are not always satisfied, particularly under strong magnetic fields where $l_{B}$ approaches the lattice scale, as in moiré systems, or in materials with relatively large node separations. Similar effects can also arise in two and three dimensional Dirac materials [Wehling2014] with two or more Dirac cones in the spectrum, e.g. graphene [Castro_2009, Katsnelson2007], and surfaces of three dimensional topological insulators [Hasan2010, Qi2011]. In such cases, the low-energy continuum description breaks down, failing to capture crucial lattice-specific phenomena.
@@
The paper is organized as follows. In Section II we present a summary of our main results. In Section III, we revisit the low-energy continuum description of gap opening induced by an orbital magnetic field. As we mentioned before, Ref. Saykin_Rodinov_2018 found the gap to be monotonic with the field, whereas Ref. Devakul_Parameswaran_2021 found the gap can exhibit periodic oscillations. In this section, we provide a unified description of the phenomena to set the stage for the richer phenomenology in a lattice description. The analysis in the presence of a lattice is carried out in Section IV, which contains the main results of this work. We conclude in Section V.
@@ II Model and Summary of the results
(i) For $\gamma<0$, the system displays three phases as a function of the separation between the Weyl nodes, $Q$, see Fig. 2(a3). For small $Q$, the system is a normal insulator (NI), whereas for large $Q$ it is a layered Chern insulator (LCI). We denote the critical momentum space separation at which the transition occurs by $Q^{*}$. These two insulating phases are separated by a gapless WSM phase. The width (range of $Q$ values) of the WSM region is exponentially small in $\phi_{B}^{-1}$. This is associated with the finite bandwidth of the LLs which is exponentially small in $\phi_{B}^{-1}$. For fixed $Q$ inside the insulating phases, the gap increases monotonically with the field.
@@ III Continuum description
In this Section, we present a unified description of the low-energy continuum behavior of our model, previously discussed in Refs. [Chan_Lee_2017, Saykin_Rodinov_2018, Devakul_Parameswaran_2021].
@@
which depends on the anisotropy of the Weyl cone. The above condition was derived before in Ref. Chan_Lee_2017.
@@ III.1 Low energy continuum model
As we have already mentioned, the anisotropy parameter $\gamma$ (defined in Eq. 2) controls the structure of the gap as a function of $Ql_{B}$. For an arbitrary $t$, the problem of computing the gap analytically becomes challenging because one would require solving a fourth-order differential equation. However, as shown in Ref. Devakul_Parameswaran_2021, a wealth of information on how the induced gap behaves as a function of magnetic flux or WN separation can be extracted analytically without solving the full eigenvalue problem.
@@ III.2 Zero energy solutions
Following Ref. Devakul_Parameswaran_2021 it is possible to obtain analytic conditions for the system to remain gapless, even in the presence of a magnetic field. The authors in Ref. Devakul_Parameswaran_2021 considered the case of $t=1$. Here we generalize the treatment to arbitrary $t\neq 0$ (the case $t=0$ is discussed in detail below).
@@
where $h=\left[a^{\dagger}a-\delta+\eta(a-a^{\dagger})\right]$ , $\delta=\frac{k_{0}^{2}l_{B}^{2}}{2t}$, $\eta=\frac{k_{0}l_{B}}{\sqrt{2t^{3}}}\frac{v_{y}}{v_{x}}$, and $\omega=\frac{v_{x}t}{k_{0}l_{B}^{2}}$. Note that $H_{\phi}^{2}=\omega^{2}\left(\begin{array}[]{c c}h^{\dagger}h&0\\ 0&hh^{\dagger}\end{array}\right)$. Hence, $H_{\phi}$ has a zero-energy eigenvalue if and only if $h$ has a zero eigenvalue. The eigenvalues of $h$ are
@@ III.3 Structure of the induced gap
We can understand the monotonic regime from the special case $t=0$, which was studied in Ref. [Saykin_Rodinov_2018]. Note that for $t=0$, the anisotropy parameter $\gamma$ is always negative for any finite value of the ratio $v_{y}/v_{x}$. For $t=0$, the problem of finding the spectrum in a uniform magnetic field along $z$ can be mapped to a Schrodinger equation with an asymmetric double-well potential. The authors in Ref. [Saykin_Rodinov_2018] found that the induced gap is nonperturbative in magnetic field. They computed the energy gap ($\Delta$) using the WKB approximation and found it to be exponentially small in $Ql_{B}$:
@@ III.4 Discussion
Several comments are in order. The oscillation and the periodic closing of the energy gap cannot be understood from the Onsager theory; its origin is purely a quantum phenomenon. The Fermi surface at zero energy consists only of two isolated points- the two Weyl nodes. At energies away from the energy of WNs, the Fermi surface consists of two disjoint and identical Fermi pockets around the Weyl nodes in the $k_{x}$-$k_{y}$ plane. When an orbital magnetic field is applied it induces a quantum tunneling from one Fermi pocket to the other, a phenomenon known as magnetic breakdown [Cohen_1961, Kaganov_1983, Shoenberg_1984]. It is this quantum tunneling between the two Fermi pockets around the WNs which leads to a gap opening in the spectrum. The probability of tunneling is exponentially small in the distance between the Fermi pockets and the height of the barrier, giving rise to the exponential decay in $(Ql_{B})^{2}$. For $\gamma>0$ the Fermi pockets are in crescent shape with two close encounters as shown in Fig. 2(b2). This gives two possible tunneling paths for the magnetic breakdown. The interference between these gives rise to the periodic gap closings. The area enclosed by the two paths is $A$ as defined in Eq. (7) accounting for the period of oscillations. Note that this picture relies on mirror symmetry which ensures that the two tunneling paths have exactly the same amplitude. If the symmetry were absent, the destructive interference would be imperfect, and the periodic gap closings would not occur. On the other hand for $\gamma<0$ the Fermi pockets have an elliptical shape as shown in Fig. 2(a2) with only one close encounter between them, leading to a monotonic behavior of the gap.
@@ IV Topological nature of the gapped phases on a lattice
The lattice model allows us to discuss the topological nature of the gapped phases induced by orbital magnetic fields. It is instructive to start from the Weyl semimetal in the absence of a magnetic field. Then, one can introduce the $k_{x}$-dependent Chern number $C(k_{x})$, which measures the Chern number in the $k_{y}$-$k_{z}$ plane at fixed $k_{x}$. For the lattice model defined in Eq. 1, $C(k_{x})=1$ for all values of $k_{x}$ in between the two Weyl nodes, $k_{0}<k_{x}<k_{0}$, and it is zero otherwise. For a system in a slab geometry with a surface in the $xy$ plane, this gives rise to Fermi arc states on the surface connecting the projection of the two Weyl nodes on the $k_{x}$-$k_{y}$ surface BZ.
@@
We now consider the effect of a magnetic field. When the two WNs are close to the center of the BZ, the Fermi arcs are short (see Fig. 4(a)). We then intuitively expect the WNs to gap out due to $Q$ tunneling (tunneling within the first BZ), and the Fermi arcs to disappear in the presence of a magnetic field. This corresponds to a topologically trivial insulating phase, which we refer to as normal insulator (NI).
@@
Figure 4: Energy spectrum (along $k_{x}$) in a slab geometry along $z$ direction with thickness $L_{z}=50$ in units of lattice constant for $\gamma\approx-0.30$ (negative). Energies are given in units of $v_{x}$. Upper panels show the spectrum for zero field, highlighting the zero energy Fermi-arc surface states (a) for small ($Q=1$) and (b) large ($Q=5$) separation between the Weyl nodes. Bottom panels show the spectrum in presence of an external field applied along $z$ direction corresponding to a flux of $\phi_{B}/\phi_{0}=1/q$ with $q=10$. (c) Fermi-arc states get fully gapped out (NI phase). (d) Fermi-arc states (highlighted in red) are extended to the edge of the magnetic BZ along $k_{x}$ (LCI phase).
@@
On the other hand, when the Weyl nodes approach each other across the edges of the BZ, the Fermi arcs are long, nearly spanning the full BZ in the $k_{x}$ direction (see Fig. 4(b)). Now, when we apply a magnetic field, we expect the Weyl nodes to gap out due to $Q^{\prime}$ tunneling (tunneling across the edge of the BZ), but the gapless Fermi arc states to survive. This corresponds to the LCI phase.
@@
Figure 5: Behavior of the gap for $\gamma<0$. (a) Zero field energy spectrum is showing two Weyl cones separated along $k_{x}$. (b) In presence of an orbital $^{\bf B}\parallel\hat{z}$, Weyl cones are gapped out. Only two Landau bands near zero energy are shown. For flux $\phi_{B}\sim\phi_{0}$, the bands have a finite bandwidth ($W$). (c) Energy gap ($\Delta$) as a function of Weyl nodes’ separation for flux values $\phi_{B}/\phi_{0}=1/q=1/10,1/20,1/50$. Magnetic field immediately creates a gap in the spectrum for almost any value of separation between the Weyl nodes. The gap is exponentially small $\sim\exp(-Q^{2}l_{B}^{2})$. For intermediate values of separation, WSM state survives but in an exponentially small vicinity of $k_{0}$ (seen most clearly for $\phi_{B}=1/3$). (d) The width of the gapless region (denoted by $\Delta Q$ in (c)) where the system remains in the WSM phase falls exponentially with $\phi_{B}^{-1}$. The bandwidth (denoted as $W$ in (b)) also falls as $\exp(-\phi_{B}^{-1})$.
@@
We note that this behavior is expected independently of the sign of $\gamma$. These notions can be made more concrete by introducing the magnetic BZ, which we discuss next. In the magnetic BZ, the momenta are good quantum numbers, making it possible to track the evolution of the Weyl nodes and Fermi arcs in the presence of a magnetic field. Figure 4(c,d) shows the spectrum on the magnetic BZ for the $Q$ and $Q^{\prime}$ tunneling cases, respectively, where the NI and LCI phases are observed. These results will be discussed in more detail below. Interestingly, the more careful analysis will also reveal the appearance of an additional gapped insulating phase, LCI, which alternates with the NI phase in the case $\gamma>0$.

### v2

@@ I Introduction
Weyl semimetals (WSMs) [1, 2, 3, 4, 5, 6, 7] represent a robust gapless topological phase of quantum matter in three dimensions. As long as spatial translation symmetry is maintained, the only way to eliminate this phase is by bringing pairs of Weyl nodes (WNs) with opposite chiralities together, allowing them to annihilate each other. WSMs exhibit a variety of exotic phenomena – the chiral anomaly [8, 9, 10, 11], chiral anomaly-induced negative magnetoresistance [12, 13, 14, 15, 16, 17, 18, 19], planar Hall effect [20, 21, 22, 23, 24], Fermi-arc-mediated quantum oscillations, and three-dimensional quantum Hall effect [25, 26, 27, 28, 29, 30, 31, 32, 33, 34] – all of which typically rely on the presence of an external orbital magnetic field. However, when the magnetic field strength is sufficiently high, it can couple WNs of opposite chiralities and open a finite gap in the spectrum, thereby destabilizing the WSM state [35, 36, 37, 38, 39, 40, 41, 42, 43].
@@
An external magnetic field that is not aligned with the separation direction of WNs can couple nodes of opposite chirality, resulting in the opening of a gap in the bulk electronic spectrum. This phenomenon has been the subject of several theoretical investigations based on low-energy continuum models. Notably, Refs. 40, 41 demonstrated that the application of an orbital magnetic field immediately induces a gap, which is exponentially small, scaling as $\sim\exp[-(Ql_{B})^{2}]$, where $Q$ represents the momentum-space separation between the Weyl nodes and $l_{B}=\sqrt{\hbar/(eB)}$ is the magnetic length associated with the applied field. In these models, the gap increases monotonically as the magnetic field strength is raised. Furthermore, Ref. 44 revealed that in systems where the Weyl fermion velocities exhibit anisotropy, the induced gap can display oscillatory behavior as a function of the magnetic field strength, adding further complexity to the field-induced modification of the electronic structure.
@@
It is important to note that these results, derived from continuum approximations, are strictly valid within a specific regime of parameters: when the separation between the Weyl nodes ($Q$) is much smaller than the linear dimensions of the Brillouin zone (BZ), and when the magnetic length $l_{B}$ is considerably larger than the lattice constant $a$, so that lattice effects can be safely ignored. However, these conditions are not always satisfied, particularly under strong magnetic fields where $l_{B}$ approaches the lattice scale, as in moiré systems, or in materials with relatively large node separations. Similar effects can also arise in two and three dimensional Dirac materials [45] with two or more Dirac cones in the spectrum, e.g. graphene [46, 47], and surfaces of three dimensional topological insulators [48, 49]. In such cases, the low-energy continuum description breaks down, failing to capture crucial lattice-specific phenomena.
@@
The paper is organized as follows. In Section II we present a summary of our main results. In Section III, we revisit the low-energy continuum description of gap opening induced by an orbital magnetic field. As we mentioned before, Ref. 40 found the gap to be monotonic with the field, whereas Ref. 44 found the gap can exhibit periodic oscillations. In this section, we provide a unified description of the phenomena to set the stage for the richer phenomenology in a lattice description. The analysis in the presence of a lattice is carried out in Section IV, which contains the main results of this work. We conclude in Section V.
@@ II Model and Summary of the results
(i) For $\gamma<0$, the system displays three phases as a function of the separation between the Weyl nodes, $Q$, see Fig. 2(a3). For small $Q$, the system is a normal insulator (NI), whereas for large $Q$ it is a layered Chern insulator (LCI). The LCI phase can be viewed as a collection of two-dimensional Chern insulators stacked along the $x$ direction, hosting chiral surface states on surfaces that result from termination along the $y$ and $z$ directions.
We denote the critical momentum space separation at which the transition occurs by $Q^{*}$. These two insulating phases are separated by a gapless WSM phase. The width (range of $Q$ values) of the WSM region is exponentially small in $\phi_{B}^{-1}$. This is associated with the finite bandwidth of the LLs which is exponentially small in $\phi_{B}^{-1}$. For fixed $Q$ inside the insulating phases, the gap increases monotonically with the field.
@@ III Continuum description
In this Section, we present a unified description of the low-energy continuum behavior of our model, previously discussed in Refs. [38, 40, 44].
@@
which depends on the anisotropy of the Weyl cone. The above condition was derived before in Ref. 38.
@@ III.1 Low energy continuum model
As we have already mentioned, the anisotropy parameter $\gamma$ (defined in Eq. 2) controls the structure of the gap as a function of $Ql_{B}$. For an arbitrary $t$, the problem of computing the gap analytically becomes challenging because one would require solving a fourth-order differential equation. However, as shown in Ref. 44, a wealth of information on how the induced gap behaves as a function of magnetic flux or WN separation can be extracted analytically without solving the full eigenvalue problem.
@@ III.2 Zero energy solutions
Following Ref. 44 it is possible to obtain analytic conditions for the system to remain gapless, even in the presence of a magnetic field. The authors in Ref. 44 considered the case of $t=1$. Here we generalize the treatment to arbitrary $t\neq 0$ (the case $t=0$ is discussed in detail below).
@@
where $h=\left[a^{\dagger}a-\delta+\eta(a-a^{\dagger})\right]$ , $\delta=\frac{k_{0}^{2}l_{B}^{2}}{2t}-\frac{1}{2}$, $\eta=\frac{k_{0}l_{B}}{\sqrt{2t^{3}}}\frac{v_{y}}{v_{x}}$, and $\omega=\frac{2v_{x}t}{k_{0}l_{B}^{2}}$. Note that $H_{\phi}^{2}=\omega^{2}\left(\begin{array}[]{c c}h^{\dagger}h&0\\ 0&hh^{\dagger}\end{array}\right)$. Hence, $H_{\phi}$ has a zero-energy eigenvalue if and only if $h$ has a zero eigenvalue. The eigenvalues of $h$ are
@@ III.3 Structure of the induced gap
We can understand the monotonic regime from the special case $t=0$, which was studied in Ref. [40]. Note that for $t=0$, the anisotropy parameter $\gamma$ is always negative for any finite value of the ratio $v_{y}/v_{x}$. For $t=0$, the problem of finding the spectrum in a uniform magnetic field along $z$ can be mapped to a Schrodinger equation with an asymmetric double-well potential. The authors in Ref. [40] found that the induced gap is nonperturbative in magnetic field. They computed the energy gap ($\Delta$) using the WKB approximation and found it to be exponentially small in $Ql_{B}$:
@@ III.4 Discussion
Several comments are in order. The oscillation and the periodic closing of the energy gap cannot be understood from the Onsager theory; its origin is purely a quantum phenomenon. The Fermi surface at zero energy consists only of two isolated points- the two Weyl nodes. At energies away from the energy of WNs, the Fermi surface consists of two disjoint and identical Fermi pockets around the Weyl nodes in the $k_{x}$-$k_{y}$ plane. When an orbital magnetic field is applied it induces a quantum tunneling from one Fermi pocket to the other, a phenomenon known as magnetic breakdown [50, 51, 52]. It is this quantum tunneling between the two Fermi pockets around the WNs which leads to a gap opening in the spectrum. The probability of tunneling is exponentially small in the distance between the Fermi pockets and the height of the barrier, giving rise to the exponential decay in $(Ql_{B})^{2}$. For $\gamma>0$ the Fermi pockets are in crescent shape with two close encounters as shown in Fig. 2(b2). This gives two possible tunneling paths for the magnetic breakdown. The interference between these gives rise to the periodic gap closings. The area enclosed by the two paths is $A$ as defined in Eq. (7) accounting for the period of oscillations. Note that this picture relies on mirror symmetry which ensures that the two tunneling paths have exactly the same amplitude. If the symmetry were absent, the destructive interference would be imperfect, and the periodic gap closings would not occur. On the other hand for $\gamma<0$ the Fermi pockets have an elliptical shape as shown in Fig. 2(a2) with only one close encounter between them, leading to a monotonic behavior of the gap.
@@ IV Topological nature of the gapped phases on a lattice
The lattice model allows us to discuss the topological nature of the gapped phases induced by orbital magnetic fields. It is instructive to start from the Weyl semimetal in the absence of a magnetic field. Then, one can introduce the $k_{x}$-dependent Chern number $C(k_{x})$, which measures the Chern number in the $k_{y}$-$k_{z}$ plane at fixed $k_{x}$ of the occupied bands. For the lattice model defined in Eq. 1, $\sum_{n\in occ}C_{n}(k_{x})=1$ for all values of $k_{x}$ in between the two Weyl nodes, $k_{0}<k_{x}<k_{0}$, and it is zero otherwise. For a system in a slab geometry with a surface in the $xy$ plane, this gives rise to Fermi arc states on the surface connecting the projection of the two Weyl nodes on the $k_{x}$-$k_{y}$ surface BZ.
@@
We now consider the effect of a magnetic field. When the two WNs are close to the center of the BZ, the Fermi arcs are short (see Fig. 4(a1)). We then intuitively expect the WNs to gap out due to $Q$ tunneling (tunneling within the first BZ), and the Fermi arcs to disappear in the presence of a magnetic field [43]. This corresponds to a topologically trivial insulating phase, which we refer to as normal insulator (NI).
@@
Figure 4: Energy spectrum (along $k_{x}$) in a slab geometry along $z$ direction with thickness $L_{z}=50$ in units of lattice constant for $\gamma\approx-0.30$ (negative). Energies are given in units of $v_{x}$. Left column shows the slab spectrum for zero field, highlighting the zero energy Fermi-arc surface states (a1) for small ($Q=1$) and (b1) large ($Q=5$) separation between the Weyl nodes. Right column shows the slab spectrum in presence of an external field applied along $z$ direction corresponding to a flux of $\phi_{B}/\phi_{0}=1/q$ with $q=10$. (a2) Fermi-arc states get fully gapped out (NI phase). (b2) The Fermi-arc states of the WSM evolve to the surface states of a LCI phase (highlighted in red). The surface states in the LCI phase are extended to the edge of the magnetic BZ along $k_{x}$.
@@
On the other hand, when the Weyl nodes approach each other across the edges of the BZ, the Fermi arcs are long, nearly spanning the full BZ in the $k_{x}$ direction (see Fig. 4(b1)). Now, when we apply a magnetic field, we expect the Weyl nodes to gap out due to $Q^{\prime}$ tunneling (tunneling across the edge of the BZ), but the gapless Fermi arc states to survive [43]. This corresponds to the LCI phase.
@@
Figure 5: Behavior of the gap for $\gamma<0$. (a) Zero field energy spectrum is showing two Weyl cones separated along $k_{x}$. (b) In presence of an orbital $^{\bf B}\parallel\hat{z}$, Weyl cones are gapped out. Only two Landau bands near zero energy are shown. For flux $\phi_{B}\sim\phi_{0}$, the bands have a finite bandwidth ($W$). (c) Energy gap ($\Delta$) as a function of Weyl nodes’ separation for flux values $\phi_{B}/\phi_{0}=1/q=1/3,1/10,1/20$. Magnetic field immediately creates a gap in the spectrum for almost any value of separation between the Weyl nodes. The gap is exponentially small $\sim\exp(-Q^{2}l_{B}^{2})$. For intermediate values of separation, WSM state survives but in an exponentially small vicinity of $k_{0}$ (seen most clearly for $\phi_{B}/\phi_{0}=1/3$). (d) The width of the gapless region (denoted by $\Delta Q$ in (c)) where the system remains in the WSM phase falls exponentially with $\phi_{B}^{-1}$. The bandwidth (denoted as $W$ in (b)) also falls as $\exp(-\phi_{B}^{-1})$.
@@
We note that this behavior is expected independently of the sign of $\gamma$. These notions can be made more concrete by introducing the magnetic BZ, which we discuss next. In the magnetic BZ, the momenta are good quantum numbers, making it possible to track the evolution of the Weyl nodes and Fermi arcs in the presence of a magnetic field. Figure 4(a2,b2) shows the spectrum on the magnetic BZ for the $Q$ and $Q^{\prime}$ tunneling cases, respectively, where the NI and LCI phases are observed. These results will be discussed in more detail below. Interestingly, the more careful analysis will also reveal the appearance of an additional gapped insulating phase, LCI, which alternates with the NI phase in the case $\gamma>0$.
As shown in this section, the behavior of the Fermi-arc surface states discussed above in the presence of a magnetic field is consistent with the nonperturbative gap-opening mechanism in the bulk Weyl semimetal. Fermi-arc states with a short arc length disappear once the magnetic field is applied. In this regime, the surface states acquire a finite lifetime, as electrons initially localized at the surface can escape into the bulk through chiral Landau levels. This fate of short Fermi arcs has also been predicted in a recent continuum analysis [53]. In contrast, Fermi-arc states with a long arc length persist and smoothly evolve into the chiral surface states of the LCI phase.